# SHIELD: Suppressing Hallucinations In LVLM Encoders via Bias and VuLnerability Defense

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) excel in diverse cross-modal tasks. However, object hallucination, where models produce plausible but inaccurate object descriptions, remains a significant challenge. In contrast to previous work focusing on LLM components, this paper is the first to trace LVLM hallucinations to visual encoders and identifies three key issues: statistical bias, inherent bias, and vulnerability. To address these challenges, we propose SHIELD, a training-free framework that mitigates hallucinations through three strategies: re-weighting visual tokens to reduce statistical bias, introducing noise-derived tokens to counter inherent bias, and applying adversarial attacks with contrastive decoding to address vulnerability. Experiments demonstrate that SHIELD effectively mitigates object hallucinations across diverse benchmarks and LVLM families. Moreover, SHIELD achieves strong performance on the general LVLM benchmark, highlighting its broad applicability. Code will be released.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) [2, 23, 8] combine visual and textual information and have advanced significantly in cross-modal tasks. Despite these advances, they suffer from object hallucination, generating object descriptions that may appear reasonable but misrepresent the image, either by misidentifying attributes (e.g., color, quantity, position) or by introducing non-existent objects. This issue poses reliability and safety risks in domains such as healthcare [12, 35], autonomous systems [5, 37], and robotics [24, 21].
@@
Various approaches have been proposed to mitigate object hallucinations. Early efforts, such as fine-grained modality alignment [3] and data augmentation to reduce co-occurrence bias [16, 30], were designed for small-scale VLMs but fail to generalize to LVLMs [15, 36]. More recent research falls into two categories: training-required and training-free methods. Training-required methods, including preference optimization [26], post-hoc revisers [39], and RLHF [31], improve factual consistency but demand substantial human and computational resources. In contrast, training-free methods, such as contrasting outputs from distorted inputs [17] or applying over-trust penalties during decoding [13], offer a more efficient alternative. However, these approaches primarily focus on the LLM component, leaving the role of visual encoders underexplored.
@@
Experiments demonstrate that SHIELD consistently improves performance on object hallucination benchmarks, including CHAIR [30], POPE [19], the hallucination subset of MME [10], and GPT4o-aided evaluations on LLaVA-Bench [23]. Moreover, these improvements are observed across diverse LVLM families, such as LLaVA [23], InstructBLIP [8], and Qwen-VL [2]. Beyond object hallucination mitigation, SHIELD also enhances general perception capabilities, as evidenced by improvements on the full MME benchmark [10], highlighting its broader applicability.
@@ 2.1 Large Vision Language Models (LVLMs)
Recent advances in large-scale foundation models and multimodal learning have accelerated the development of Large Vision-Language Models (LVLMs). By combining Large Language Models (LLMs) [1, 4, 6, 7, 11, 29, 32, 33, 34] with cross-modal frameworks such as CLIP [28] and BLIP [18], LVLMs integrate visual and textual information for more comprehensive understanding. Nevertheless, LVLMs across different architectures, including LLaVA-1.5 [22], InstructBLIP [8], and Qwen-VL [2], still suffer from hallucinations, particularly in fine-grained object recognition and challenging visual grounding. Such errors, often involving non-existent objects or misidentified attributes, remain a key challenge for reliability in real-world applications.
@@ 2.2 Object Hallucination in LVLMs
Approaches to mitigating object hallucinations in LVLMs can be grouped into training-required and training-free methods. Training-required methods reduce hallucinations by optimizing model parameters or training auxiliary modules. Prominent methods include CLIP-DPO [26], leveraging CLIP-based similarity ranking for preference optimization; LURE [39], using post-hoc revisers to align text with visual input; and LLaVA-RLHF [31], incorporating human feedback through reinforcement learning. Training-free methods improve decoding without modifying model. Representative approaches include Visual Contrastive Decoding (VCD) [17], which contrasts outputs from natural and blur inputs to mitigate hallucinations, and OPERA [13], which reduces overconfidence through penalty mechanisms and token-level adjustments. While effective, these methods rarely address the bias and vulnerability of visual encoders, which this work aims to address.
@@ 3.1.1 Statistical and Inherent Bias in Visual Encoder
Most LVLMs adopt visual encoders derived from pretrained CLIP models. Although these encoders benefit from large-scale pretraining, they are influenced by the imbalanced distribution of visual concepts in the pretraining data. Specifically, certain visual concepts appear far more frequently than others, while rare or context-dependent elements are severely underrepresented [27]. As a result, the model develops a strong inductive bias toward frequent patterns and dominant objects, giving rise to both statistical and inherent bias.
Statistical bias denotes the visual encoder’s over-reliance on frequent visual patterns in the pretraining data, causing overemphasis on the corresponding tokens with disproportionately high L2 activation values [9]. This overemphasis distorts the downstream LLM’s perception of fine-grained details by directing attention to overweighted tokens [14], often resulting in hallucinations. Analysis of LLaVA-1.5’s responses and visual tokens on the POPE COCO subset (Figure 2(a)) shows that the proportion of hallucinated samples grows with stronger token overemphasis, measured by the peak-to-average L2 ratio (the deviation of the highest-norm token from the mean among visual tokens).
@@ 3.1.2 Vulnerability in Visual Encoder
The vulnerability of visual encoders is another key factor contributing to object hallucinations. It arises from their limited robustness to noise and subtle perturbations [25], making them susceptible to constructing inaccurate visual representations under such disturbances. As shown in Figure 2(c), the performance of LLaVA-1.5 drops sharply on the POPE COCO subset even with a few attack steps, demonstrating that minor perturbations can exploit this weakness and yield unreliable features.
@@ 3.2.4 Address Vulnerability
Following [17], an adaptive plausibility constraint is introduced to avoid implausible outputs. Only tokens with probabilities no smaller than a fraction $\beta$ of the maximum are retained:
@@ 4.1 Implementation Details
To evaluate the effectiveness of SHIELD in mitigating hallucinations, three representative LVLMs were selected: LLaVA-1.5 [22], InstructBLIP [8], and Qwen-VL [2]. SHIELD was compared against the corresponding vanilla LVLMs and two recent training-free methods, VCD [17] and OPERA [13]. Following their original setups, vanilla LVLMs and VCD adopted sampling-based decoding, while OPERA employed beam search decoding with a penalty term on logits to reduce overconfidence. For SHIELD, sampling-based decoding was used, drawing from the modified post-softmax distribution. Unless otherwise specified, $\alpha=2$, $\beta=0.35$, $K=32$, and $l=0.02$ were applied across all LVLMs, where $\alpha$ controls the strength of contrastive decoding, $\beta$ sets the truncation threshold in the plausibility constraint, $K$ denotes the number of noise inputs for estimating inherent bias, and $l$ is the learning rate for optimizing the attack tensor. All experiments were conducted on a single RTX A6000 GPU.
@@ 4.2 Quantitative Results
CHAIR Evaluation. The Caption Hallucination Assessment with Image Relevance (CHAIR) metric [30] assesses object hallucinations in image captioning by calculating the proportion of object references absent from the ground-truth annotations. It comprises two levels of evaluation, $C_{S}$ (sentence-level) and $C_{I}$ (instance-level):
$C_{S}=\frac{|\text{sentences with hallucinated object}|}{|\text{all sentences}|},\quad C_{I}=\frac{|\text{hallucinated objects}|}{|\text{all objects mentioned}|}.$
Following the setup in [13], we conduct evaluations on 500 randomly sampled images from the COCO 2014 validation dataset [20], using the prompt “Please describe this image in detail.” For fair comparison, all generated captions are truncated to a maximum length of 512 tokens.
As shown in Table 2, SHIELD consistently outperforms all previous training-free decoding methods on both $C_{S}$ and $C_{I}$, achieving up to 18% improvement over the second-best method, OPERA, on LLaVA-1.5. This performance gain stems from SHIELD’s ability to counter biases and vulnerability in the visual encoder, thereby reducing hallucination risk in detailed descriptions.
@@
GPT4o Assisted Evaluation. While CHAIR effectively evaluates object-level hallucinations, it fails to capture errors in attributes, locations, or relations. To complement this, we employ GPT4o, a strong multi-modal assistant, to assess LVLM outputs on the LLaVA-Bench dataset. GPT4o scores responses on correctness (C) and detailedness (D) from 0–10, with higher correctness indicating fewer hallucinations. The evaluation explicitly targets objects mentioned but absent from the image, as well as errors in attributes, colors, positions, or relationships. Further details are provided in the Appendix A.1.
@@
POPE Evaluation. Similar to CHAIR, the Polling-based Object Probing Evaluation (POPE) [19] assesses existence-level hallucinations in LVLMs. It adopts a VQA-style format (e.g., “Is there a {object} in the image?”) to test whether models correctly associate images with specific objects. POPE includes three splits: “random” for random objects, “popular” for frequent objects, and “adversarial” for objects semantically related to those in the image. The evaluation is conducted on three subsets: COCO, A-OKVQA, and GQA. Additional results are provided in Appendix B.1.
@@
MME Hallucination Subset Evaluation. Although POPE adopts a VQA format effective for evaluating object-existence-level hallucinations, it does not capture attribute-level aspects such as count, position, and color. To address this limitation, the MME hallucination subsets [10] provide a more comprehensive benchmark. Following [38], we evaluate object-level hallucinations using the existence and count subsets, and attribute-level hallucinations using the position and color subsets. Performance is reported using the combined metrics of accuracy and accuracy+ as defined in the official implementation.
LVLM General Evaluation. To evaluate the overall performance of SHIELD-enhanced LVLMs, we conduct experiments on the full MME benchmark [10] using the LLaVA-1.5 7B model. The benchmark covers ten perception-related subtasks (including four hallucination-related) and four cognition-oriented ones. Performance is reported using both accuracy and accuracy+ as defined in the official implementation.
@@ Appendix A Detailed Experimental Setup
A.1 GPT4o Assisted Evaluation
Following [38], we use GPT4o to evaluate the vanilla LVLM, VCD, OPERA, and our proposed SHIELD model. For each LVLM and image, we generate a description using the prompt “Please describe this image in detail.” Using the evaluation prompt in Table 7, GPT4o scores the four descriptions from 0 to 10 on two aspects: correctness, which measures how well the description matches the image, giving higher scores to accurate descriptions and lower scores to those with hallucinated content, and detailedness, which assesses how fully the description captures image details. The prompt instructs GPT4o to ignore bias from the order of inputs and focus on inconsistencies such as objects mentioned but not present, or incorrect colors, positions, and relationships.

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) excel in diverse cross-modal tasks. However, object hallucination, where models produce plausible but inaccurate object descriptions, remains a significant challenge. In contrast to previous work focusing on LLM components, this paper is the first to trace LVLM hallucinations to visual encoders and identifies three key issues: statistical bias, inherent bias, and vulnerability. To address these challenges, we propose SHIELD, a training-free framework that mitigates hallucinations through three strategies: re-weighting visual tokens to reduce statistical bias, introducing noise-derived tokens to counter inherent bias, and applying adversarial attacks with contrastive decoding to address vulnerability. Experiments demonstrate that SHIELD effectively mitigates object hallucinations across diverse benchmarks and LVLM families. Moreover, SHIELD achieves strong performance on the general LVLM benchmark, highlighting its broad applicability. Code is available at https://github.com/hukcc/SHIELD.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) (Bai et al., 2023b; Liu et al., 2023b; Dai et al., 2023) combine visual and textual information and have advanced significantly in cross-modal tasks. Despite these advances, they suffer from object hallucination, generating object descriptions that may appear reasonable but misrepresent the image, either by misidentifying attributes (e.g., color, quantity, position) or by introducing non-existent objects. This issue poses reliability and safety risks in domains such as healthcare (Hu et al., 2023; Wang et al., 2023b), autonomous systems (Chen et al., 2024a; Wu et al., 2023), and robotics (Mai et al., 2023; Liu et al., 2023a).
@@
Various approaches have been proposed to mitigate object hallucinations. Early efforts, such as fine-grained modality alignment (Biten et al., 2022) and data augmentation to reduce co-occurrence bias (Kim et al., 2023; Rohrbach et al., 2018), were designed for small-scale VLMs but fail to generalize to LVLMs (Kaplan et al., 2020; Wei et al., 2022). More recent research falls into two categories: training-required and training-free methods. Training-required methods, including preference optimization (Ouali et al., 2024), post-hoc revisers (Zhou et al., 2024), and RLHF (Sun et al., 2024a), improve factual consistency but demand substantial human and computational resources. In contrast, training-free methods, such as contrasting outputs from distorted inputs (Leng et al., 2024) or applying over-trust penalties during decoding (Huang et al., 2024), offer a more efficient alternative. However, these approaches primarily focus on the LLM component, leaving the role of visual encoders underexplored.
@@
Experiments demonstrate that SHIELD consistently improves performance on object hallucination benchmarks, including CHAIR (Rohrbach et al., 2018), POPE (Li et al., 2023), the hallucination subset of MME (Fu et al., 2023), and GPT-4o-aided evaluations on LLaVA-Bench (Liu et al., 2023b). Moreover, these improvements are observed across diverse LVLM families, such as LLaVA (Liu et al., 2023b), InstructBLIP (Dai et al., 2023), and Qwen-VL (Bai et al., 2023b). Beyond object hallucination mitigation, SHIELD also enhances general perception capabilities, as evidenced by improvements on the full MME benchmark (Fu et al., 2023), highlighting its broader applicability.
@@ 2.1 Large Vision Language Models (LVLMs)
Recent advances in large-scale foundation models and multimodal learning have accelerated the development of Large Vision-Language Models (LVLMs). By combining Large Language Models (LLMs) (Bai et al., 2023a; Brown et al., 2020; Chiang et al., 2023; Chowdhery et al., 2023; Gilardi et al., 2023; Raffel et al., 2020; Taori et al., 2023; Tay et al., 2023; Touvron et al., 2023) with cross-modal frameworks such as CLIP (Radford et al., 2021) and BLIP (Li et al., 2022), LVLMs integrate visual and textual information for more comprehensive understanding. Nevertheless, LVLMs across different architectures, including LLaVA-1.5 (Liu et al., 2024), InstructBLIP (Dai et al., 2023), and Qwen-VL (Bai et al., 2023b), still suffer from hallucinations, particularly in fine-grained object recognition and challenging visual grounding. Such errors, often involving non-existent objects or misidentified attributes, remain a key challenge for reliability in real-world applications.
@@ 2.2 Object Hallucination in LVLMs
Approaches to mitigating object hallucinations in LVLMs can be grouped into training-required and training-free methods. Training-required methods reduce hallucinations by optimizing model parameters or training auxiliary modules. Prominent methods include CLIP-DPO (Ouali et al., 2024), leveraging CLIP-based similarity ranking for preference optimization; LURE (Zhou et al., 2024), using post-hoc revisers to align text with visual input; and LLaVA-RLHF (Sun et al., 2024a), incorporating human feedback through reinforcement learning. Training-free methods improve decoding without modifying model. Representative approaches include Visual Contrastive Decoding (VCD) (Leng et al., 2024), which contrasts outputs from natural and blur inputs to mitigate hallucinations, OPERA (Huang et al., 2024), which reduces overconfidence through penalty mechanisms and token-level adjustments, HALC (Chen et al., 2024b), which employs adaptive focal-contrast decoding to provide token-wise visual grounding during inference, MARINE (Zhao et al., 2025a), which incorporates image-grounded guidance from external vision models, and VTI (Liu et al., 2025), which steers latent representations to stabilize vision features at test time. While effective, these methods rarely address the bias and vulnerability of visual encoders, which this work aims to address.
@@ 3.1.1 Statistical and Inherent Bias in Visual Encoder
Most LVLMs adopt visual encoders derived from pretrained CLIP models. Although these encoders benefit from large-scale pretraining, they are influenced by the imbalanced distribution of visual concepts in the pretraining data. Specifically, certain visual concepts appear far more frequently than others, while rare or context-dependent elements are severely underrepresented (Parashar et al., 2024). As a result, the model develops a strong inductive bias toward frequent patterns and dominant objects, giving rise to both statistical and inherent bias.
Statistical bias denotes the visual encoder’s over-reliance on frequent visual patterns in the pretraining data, causing overemphasis on the corresponding tokens with disproportionately high L2 activation values (Darcet et al., 2024). This overemphasis distorts the downstream LLM’s perception of fine-grained details by directing attention to overweighted tokens (Kang et al., 2025), often resulting in hallucinations. Analysis of LLaVA-1.5’s responses and visual tokens on the POPE COCO subset (Figure 2(a)) shows that the proportion of hallucinated samples grows with stronger token overemphasis, measured by the peak-to-average L2 ratio (the deviation of the highest-norm token from the mean among visual tokens).
@@ 3.1.2 Vulnerability in Visual Encoder
The vulnerability of visual encoders is another key factor contributing to object hallucinations. It arises from their limited robustness to noise and subtle perturbations (Mao et al., 2023), making them susceptible to constructing inaccurate visual representations under such disturbances. As shown in Figure 2(c), the performance of LLaVA-1.5 drops sharply on the POPE COCO subset even with a few attack steps, demonstrating that minor perturbations can exploit this weakness and yield unreliable features.
@@ 3.2.4 Address Vulnerability
Following (Leng et al., 2024), an adaptive plausibility constraint is introduced to avoid implausible outputs. Only tokens with probabilities no smaller than a fraction $\beta$ of the maximum are retained:
@@ 4.1 Implementation Details
To evaluate the effectiveness of SHIELD in mitigating hallucinations, three representative LVLMs were selected: LLaVA-1.5 (Liu et al., 2024), InstructBLIP (Dai et al., 2023), and Qwen-VL (Bai et al., 2023b). SHIELD was compared against the corresponding vanilla LVLMs and two recent training-free methods, VCD (Leng et al., 2024) and OPERA (Huang et al., 2024). Following their original setups, vanilla LVLMs and VCD adopted sampling-based decoding, while OPERA employed beam search decoding with a penalty term on logits to reduce overconfidence. For SHIELD, sampling-based decoding was used, drawing from the modified post-softmax distribution. Unless otherwise specified, $\alpha=2$, $\beta=0.35$, $K=32$, and $l=0.02$ were applied across all LVLMs, where $\alpha$ controls the strength of contrastive decoding, $\beta$ sets the truncation threshold in the plausibility constraint, $K$ denotes the number of noise inputs for estimating inherent bias, and $l$ is the learning rate for optimizing the attack tensor. Hyper-parameters ablation are provided in the Appendix C. All experiments were conducted on a single RTX A6000 GPU.
@@ 4.2 Quantitative Results
CHAIR Evaluation. The CHAIR metric (Rohrbach et al., 2018) measures object hallucination in image captioning by calculating the proportion of objects mentioned in captions that are not present in the ground-truth label set. It consists of two dimensions: a sentence-level score $C_{S}$ and an instance-level score $C_{I}$, which are defined as:
$C_{S}=\frac{|\text{sentences with hallucinated objects}|}{|\text{all sentences}|},\quad C_{I}=\frac{|\text{hallucinated objects}|}{|\text{all objects mentioned}|}.$
Following (Huang et al., 2024), evaluation is performed on 500 randomly selected images from the COCO 2014 validation set (Lin et al., 2014), using the prompt “Please describe this image in detail.” To ensure fairness, generated captions are truncated to a maximum of 512 tokens.
As shown in Table 2, SHIELD consistently outperforms all previous training-free methods on both $C_{S}$ and $C_{I}$, achieving up to 18% improvement over the second-best method, OPERA, on LLaVA-1.5. This performance gain stems from SHIELD’s ability to counter biases and vulnerability in the visual encoder, thereby reducing hallucination risk in detailed descriptions.
@@
GPT-4o Assisted Evaluation. While CHAIR effectively evaluates object-level hallucinations, it fails to capture errors in attributes, locations, or relations. To complement this, we employ GPT-4o, a strong multi-modal assistant, to assess LVLM outputs on the LLaVA-Bench dataset. GPT-4o scores responses on correctness (C) and detailedness (D) from 0–10, with higher correctness indicating fewer hallucinations. The evaluation explicitly targets objects mentioned but absent from the image, as well as errors in attributes, colors, positions, or relationships. Further details are provided in the Appendix A.1.
@@
POPE Evaluation. Similar to CHAIR, POPE (Li et al., 2023) evaluates existence-level hallucinations in LVLMs. It adopts a VQA-style format (e.g., “Is there a {object} in the image?”) to test whether models correctly associate images with specific objects. POPE includes three splits: “random” for random objects, “popular” for frequent objects, and “adversarial” for objects semantically related to those in the image. The evaluation is conducted on three subsets: COCO, A-OKVQA, and GQA. Additional results are provided in Appendix B.1.
MME Hallucination Subset Evaluation. Although POPE adopts a VQA format effective for evaluating object-existence-level hallucinations, it does not capture attribute-level aspects such as count, position, and color. To address this limitation, the MME hallucination subsets (Fu et al., 2023) provide a more comprehensive benchmark. Following (Yin et al., 2023), we evaluate object-level hallucinations using the existence and count subsets, and attribute-level hallucinations using the position and color subsets. Performance is reported using the combined metrics of accuracy and accuracy+ as defined in the official implementation.
Table 5: AMBER Hallucination Evaluation
Method Generative Task Discriminative Task AMBER Score
[…]
Vanilla 9.2 41.3 29.2 3.7 65.7 83.2 64.7 73.2 82.0
VCD 8.1 44.2 28.6 3.1 68.3 85.8 65.2 74.0 82.9
OPERA 8.3 43.1 31.2 2.9 76.0 79.2 83.8 81.4 86.5
[…]
Attribute and Relation Level Hallucination Evaluation. To further evaluate hallucinations at the attribute and relation level, we conduct experiments on the AMBER benchmark (Wang et al., 2023a), comparing against a series of baselines based on LLaVA-1.5 7B. AMBER offers a comprehensive suite of generative and discriminative metrics, explicitly covering existence, attribute, and relation hallucinations, making it a strong benchmark for assessing multimodal hallucination across diverse tasks. As shown in Table 5, our method achieves the highest Amber Score (88.0) and consistently outperforms all baselines across most metrics, demonstrating effective mitigation of hallucinations.
@@
LVLM General Evaluation. To evaluate the overall performance of SHIELD-enhanced LVLMs, we conduct experiments on the full MME benchmark (Fu et al., 2023) using the LLaVA-1.5 7B model. The benchmark covers ten perception-related subtasks (including four hallucination-related) and four cognition-oriented ones. Performance is reported using both accuracy and accuracy+ as defined in the official implementation.
@@ 4.3 Ablation Study
Table 8: Efficiency Comparison on CHAIR
[…]
Vanilla 48.8 2.59 15.69GB
VCD 46.8 4.89 16.52GB
OPERA 44.6 24.01 34.88GB
[…]
Table 9: Module Efficiency on CHAIR
[…]
Vanilla 2.59 15.69GB
w/ Mitigate Statistical Bias 4.64 16.56GB
w/ Reduce Inherent Bias 2.63 16.50GB
w/ Address Vulnerability 7.30 18.17GB
[…]
Inference Overhead Analysis. To assess the computational cost of SHIELD, we evaluate its hallucination ratio, runtime, and peak GPU memory usage on the CHAIR benchmark with LLaVA-1.5 7B. As shown in Table 9, SHIELD achieves the lowest hallucination ratio ($C_{S}$ = 36.6) while maintaining acceptable inference overhead. In particular, it runs substantially faster than OPERA and provides a more favorable trade-off between hallucination reduction and computational efficiency compared to VCD. To further analyze the sources of computational cost, we conduct a module-wise ablation study using the CHAIR benchmark and LLaVA-1.5 7B, as shown in Table 9. The majority of overhead arises from the Address Vulnerability and Mitigate Statistical Bias modules, both of which require caption generation. In particular, Address Vulnerability further involves adversarial tensor computation, increasing runtime to 7.30 seconds per sample and memory usage to 18.17 GB. The Mitigate Statistical Bias module, which only relies on caption generation, increases runtime to 4.64 seconds and memory to 16.56 GB. For reference, caption generation alone takes 2.05 seconds and 16.56 GB. In contrast, the Reduce Inherent Bias module introduces minimal overhead, adding only 0.04 seconds and 0.81 GB. Notably, the full SHIELD configuration incurs negligible additional cost beyond Address Vulnerability. These results indicate that most overhead is concentrated in caption generation and adversarial tensor computation, and that the overall trade-off can be flexibly adjusted by tuning caption length or the number of adversarial optimization steps.
@@ 5 Conclusion
Future work could explore the design of more robust visual encoders to further reduce bias and vulnerability, as well as the optimization of adversarial attack mechanisms to lower computational costs for real-time applications.
Ethics Statement
In this paper, we propose SHIELD, a method that mitigates bias and vulnerability to reduce hallucinations in LVLMs, thereby enhancing their safety and reliability for the community. Outputs of SHIELD may occasionally contain inappropriate content inherited from the base model, and such outputs do not reflect the authors’ views. We strictly adhere to the ICLR ethical research standards and applicable laws. To the best of our knowledge, this work complies with the General Ethical Principles.
Reproducibility Statement
We follow the ICLR reproducibility standards and ensure the reproducibility of our work. All datasets used for inference and evaluation are publicly available, and detailed experimental settings, including hyper-parameters and implementation steps, are documented in the paper and Appendix. Furthermore, we will release our code upon acceptance, enabling other researchers and practitioners to easily reproduce and extend our results.
@@ Appendix A Detailed Experimental Setup
A.1 GPT-4o Assisted Evaluation
Following (Yin et al., 2023), we use GPT-4o to evaluate vanilla LVLM, VCD, OPERA, and our proposed SHIELD. Given a LVLM and an image, descriptions are generated using the prompt, “Please describe this image in detail”. Using the evaluation prompt shown in Table 10, GPT-4o rates the four descriptions on a scale of 0 to 10 across two aspects: Correctness, which measures the consistency between the description and the image, assigning higher scores to highly consistent descriptions and lower scores to those with hallucinations, and Detailedness, which evaluates how comprehensively the description captures image details. The prompt instructs GPT-4o to disregard biases from sequential order and focus on inconsistencies, such as objects mentioned but absent in the image, including incorrect colors, positions, or relationships. Leveraging its analytical capabilities, GPT-4o performs a thorough and detailed evaluation.